OpenStudy (anonymous):

x/y=(1/y)+1 make x as a subject

OpenStudy (anonymous):

What do you mean by subject here?

OpenStudy (anonymous):

x=

OpenStudy (zarkon):

multiply both sides by y

OpenStudy (anonymous):

[x/y=1/y+1]y

OpenStudy (zarkon):

\[\frac{x}{y}=\frac{1}{y}+1\] \[y\frac{x}{y}=y\left(\frac{1}{y}+1\right)\] \[x=y\left(\frac{1}{y}+1\right)\]

OpenStudy (anonymous):

if 1/x=2+(1/y) how?

OpenStudy (zarkon):

flip both sides \[x=\frac{1}{2+\frac{1}{y}}\]